# FlexSP: Accelerating Large Language Model Training via Flexible Sequence Parallelism

### v2

@@ Abstract.
Distributed LLM Training; Sequence Parallelism; Flexible Strategies; Heterogeneous Workloads
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3676641.3715998isbn: 979-8-4007-1079-7/2025/03ccs: Computer systems organization Cloud computingccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Parallel computing methodologies
@@ 4.1.2. Cost Estimation
Furthermore, when combining sequence parallelism with ZeRO (especially ZeRO-3), we also estimate the overhead of parameter gathering and gradient synchronization, and also consider the overlapping of computation and communication like previous works (Miao et al., 2022; Wang et al., 2024). As ZeRO is orthogonal to our work, and its overhead is unrelated with the sequence parallelism nor the sequence lengths, we omit such details in this paper for clarity. Experiments show that the overall cost estimation error is below 6%, detailed in Appendix C (mya, 2025).
@@ 4.2. Sequence Blaster
This takeaway focuses on striking a balance in memory consumption across micro-batches, which is proportional to the total token number. It is designed to prevent potential out-of-memory (OOM) situations when splitting micro-batches and also to avoid under-utilization of device memory. This guidance also contributes to takeaway #1, as imbalanced token blasting leads to more micro-batches. Therefore, we design a memory-balanced micro-batch chunking algorithm based on dynamic programming, detailed in Appendix A (mya, 2025).
@@ 6.1. Experimental Setups
We conduct experiments on GPT-series LLMs of three different sizes, GPT-7B, GPT-13B, GPT-30B. Refer to Appendix B.1 (mya, 2025) for more details. We choose three different datasets, including GitHub, CommonCrawl, and Wikipedia. Fig. 2 displays the distribution of sequence lengths of these datasets. We also evaluate each system on these datasets under different maximum context length limits, i.e., 384K and 192K. The sequences exceed the maximum context length limit will be eliminated during training.
@@
We apply sequence packing for all systems. Specifically, for baseline systems Megatron-LM, DeepSpeed and FlexSP-BatchAda, we use the Best-fit Packing (Ding et al., 2024a) as introduced in §2. For FlexSP, the solver will automatically determine the sequence packing. As for the parallelism strategy, we manually tune the most efficient strategy for baseline systems under different workloads, including parallelism degrees of DP, TP, PP, CP, SP. We also apply activation checkpointing strategies for each system to accommodate model training with a context length of 384K. We fix the global batch size of each training step as 512 for all workloads, and record the average iteration time over 40 iterations after 10-iteration’s warm-up. Refer to Appendix B.2 (mya, 2025) for details.
@@ 6.6. Complexity and Scalability of FlexSP Solver
Due to the space limitation, we put more results and analysis in the Appendix (mya, 2025), including more details of our experimental setups in Appendix B, estimation accuracy of cost models in Appendix C, performance analysis of SOTA systems in Appendix D, discussion about integrating context parallelism in Appendix E. Please refer to our Appendix for more details.
@@ 7. Related Work
Efforts to optimize long context training have led to various elaborate parallel strategies, such as ring attention for LLMs (Li et al., 2023b; Liu et al., 2023; Li et al., 2024; Brandon et al., 2023), though they often suffer from communication overhead and inefficiencies with severe communication cost. These methods are orthogonal, and can be integrated into FlexSP seamlessly, detailed in Appendix E (mya, 2025).Other works aim to support long context training by extending attention mechanism (Dai et al., 2019; Beltagy et al., 2020) and optimizing position embedding  (Ding et al., 2024b; Zhu et al., 2024), which are also orthogonal.

### v3

@@ 4.1.2. Cost Estimation
Furthermore, when combining sequence parallelism with ZeRO (especially ZeRO-3), we also estimate the overhead of parameter gathering and gradient synchronization, and also consider the overlapping of computation and communication like previous works (Miao et al., 2022; Wang et al., 2024). As ZeRO is orthogonal to our work, and its overhead is unrelated with the sequence parallelism nor the sequence lengths, we omit such details in this paper for clarity. Experiments show that the overall cost estimation error is below 6%, detailed in Appendix C.
@@ 4.2. Sequence Blaster
This takeaway focuses on striking a balance in memory consumption across micro-batches, which is proportional to the total token number. It is designed to prevent potential out-of-memory (OOM) situations when splitting micro-batches and also to avoid under-utilization of device memory. This guidance also contributes to takeaway #1, as imbalanced token blasting leads to more micro-batches. Therefore, we design a memory-balanced micro-batch chunking algorithm based on dynamic programming, detailed in Appendix A.
@@ 6.1. Experimental Setups
We conduct experiments on GPT-series LLMs of three different sizes, GPT-7B, GPT-13B, GPT-30B. Refer to Appendix B.1 for more details. We choose three different datasets, including GitHub, CommonCrawl, and Wikipedia. Fig. 2 displays the distribution of sequence lengths of these datasets. We also evaluate each system on these datasets under different maximum context length limits, i.e., 384K and 192K. The sequences exceed the maximum context length limit will be eliminated during training.
@@
We apply sequence packing for all systems. Specifically, for baseline systems Megatron-LM, DeepSpeed and FlexSP-BatchAda, we use the Best-fit Packing (Ding et al., 2024a) as introduced in §2. For FlexSP, the solver will automatically determine the sequence packing. As for the parallelism strategy, we manually tune the most efficient strategy for baseline systems under different workloads, including parallelism degrees of DP, TP, PP, CP, SP. We also apply activation checkpointing strategies for each system to accommodate model training with a context length of 384K. We fix the global batch size of each training step as 512 for all workloads, and record the average iteration time over 40 iterations after 10-iteration’s warm-up. Refer to Appendix B.2 for details.
@@ 6.6. Complexity and Scalability of FlexSP Solver
Due to the space limitation, we put more results and analysis in the Appendix, including more details of our experimental setups in Appendix B, estimation accuracy of cost models in Appendix C, performance analysis of SOTA systems in Appendix D, discussion about integrating context parallelism in Appendix E. Please refer to our Appendix for more details.
@@ 7. Related Work
Efforts to optimize long context training have led to various elaborate parallel strategies, such as ring attention for LLMs (Li et al., 2023b; Liu et al., 2023; Li et al., 2024; Brandon et al., 2023), though they often suffer from communication overhead and inefficiencies with severe communication cost. These methods are orthogonal, and can be integrated into FlexSP seamlessly, detailed in Appendix E.Other works aim to support long context training by extending attention mechanism (Dai et al., 2019; Beltagy et al., 2020) and optimizing position embedding  (Ding et al., 2024b; Zhu et al., 2024), which are also orthogonal.
@@ Acknowledgements.
Appendix A Momery-balanced Micro-batch Chunking in Sequence Blaster
We illustrate the memory-balanced micro-batch chunking algorithm in Sequence Blaster based on dynamic programming. Specifically, given a batch of sequences =$\mathcal{B}=\{\mathcal{S}_{k}\}$ that has already been sorted according to takeaway #2, we split them into consecutive $M$ micro-batches, and micro-batch $\mathcal{M}_{i}$ contains all sequences $k$ satisfying $j_{i-1}\leq k<j_{i}$, where $j_{i}$ is the ending indices for $\mathcal{M}_{i}$ ($j_{0}=0$). To balance the token amount of each micro-batch, we aim to minimize the maximum total token number of each micro-batch as follows:
(23) ${\small\operatorname*{arg\,min}_{\{j_{i}\}}\max_{i\in[1,M]}\{\sum_{k\in[j_{i-1},j_{i})}s_{k}\}.}$
Again, we solve the problem via dynamic programming. Denote $DP[k][i]$ as the optimal value when blasting the first $k$ sequences into $i$ micro-batches. Starting with $DP[0][0]=0$, we can solve the problem via the following state transition formula:
(24) ${\small DP[k][i]=\min_{j\in[i-1,k-1]}\{\max\{DP[j][i-1],\sum_{l\in[j+1,k]}s_{l}\}\},}$
where $\sum_{l\in[j+1,k]}s_{l}$ represents the total token number of the $i^{th}$ micro-batch when splitting micro-batch at the $\mathcal{S}_{j}$. $DP[K][M]$ denotes the optimal solution, and the optimized values of $j_{i}$ splits the global batch data into $M$ micro-batches with balanced memory consumption.
Appendix B Details of Experimental Setups
B.1. Model Configuration
Tab.  5 presents the specific configurations of the GPT-7B, 13B and 30B models used in our experiments. The parameter number of each model in Tab. 5 is under maximum context length of 384K, where the positional embedding contains 1-2 billion parameters.
Table 5. Model configuration (384K max context length).
Model # Layers # Param Hidden Dim
GPT-7B 32 7.85B 4096
GPT-13B 40 14.03B 5120
GPT-30B 60 32.72B 6656
B.2. Protocols
For fair comparison, we manually tune the most efficient parallelism strategies for all baseline systems under different workloads. For DeepSpeed, the optimal strategy is usually among SP=64 or SP=32 with ZeRO-3, and for Megatron-LM, the optimal strategy is usually among TP=8, CP=8, or TP=16, CP=4, or TP=8, CP=4, DP=2 with ZeRO-1.
We also apply activation checkpointing strategies for each system to make sure all systems can fit the models without out-of-memory issues. For GPT-7B, activation checkpointing is unnecessary to support a 384K context length on 64 GPUs. For GPT-13B, we only checkpoint MLP layers, while for GPT-30B, almost all layers are checkpointed to support 384K context length.
Appendix C Estimation Accuracy of Cost Models
We evaluate the accuracy of the cost estimator (§4.1.2) utilized in FlexSP across diverse configurations (as shown in Tab. 1), including sequence parallelism degree, batch size, and sequence length. Fig. 9 compares the deviation of the estimated cost and the empirical execution time. As can be seen, our overhead estimator adeptly approximates the execution overhead, with discrepancies consistently remaining below 5%. The accurate estimations rendered by the estimator facilitates performance of our system.
Figure 9. Estimation accuracy.
Appendix D Performance Analysis of SOTA Systems
Here we analyze the performance of SOTA systems, i.e., DeepSpeed and Megatron-LM. As shown in Fig.  4, in most cases, DeepSpeed has similar or better performance than Megatron-LM. This is attributed to the different mechanism of CP in Megatron-LM and SP in DeepSpeed, as well as the skewness of datasets. CP usually has much more communication volumn than the All-to-All in SP. Although CP leverages the overlap between attention computation and KV transmission to hide the communication overhead, in scenarios with limited inter-node bandwidth and a majority of short sequences in datasets, the attention computation often fails to hide the communication. Therefore, Megatron-LM’s performance is usually constrained by its higher communication volume compared to DeepSpeed and FlexSP.
Appendix E Integrating Context Parallelism
Context parallelism (Li et al., 2023b; Liu et al., 2023; Li et al., 2024; Brandon et al., 2023) is also an important technique for long context training, as illustrated in §2.1.3. As a different paradigm to scatter sequences across devices, CP is orthogonal to our work, and can be integrated into FlexSP. Specifically, similar to FlexSP utilizing ZeRO and flexible SP, we can employ TP, ZeRO and CP, fix the parallelism degree of TP, and employ the flexible sequence parallelism strategy of FlexSP to achieve flexible CP, which adaptively adjusts the CP group size according to the varied-length sequences. We’ll integrate CP into our system as the future work.